OpenStudy (anonymous):

Help please with slopes!! Find the slope for the table of values below. x 1 2 3 4 y 8 12 16 20 (Points : 2) 4 8 1/8 1/4

OpenStudy (anonymous):

 Slope formula ~

OpenStudy (anonymous):

Well, a good way to think about this problem is to ask yourself: how much is y increasing for every x. for x=1, y=8 for x=2, y=12 for x=3, y=16... so you can see that everytime x runs by 1 y rises by 4 slope=rise/run=4/1

OpenStudy (anonymous):

alternatively you can you the slope formula. just plug in two coordinate points. 12-8/2-1=4/1

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

no problem. do you understand what I did?

OpenStudy (anonymous):

Yep! Thank you!

OpenStudy (anonymous):

no prob. :)